The electric potential at a point is given by
V = z(z² - 4x²). Calculate the electric field
E at that point
1
Expert's answer
2020-03-02T10:38:20-0500

"\overrightarrow{E}=-\overrightarrow{\nabla}\cdot U"


"\frac{\partial U}{\partial x}=-8zx"


"\frac{\partial U}{\partial y}=0"


"\frac{\partial U}{\partial z}=3z^2-4x^2"


"\overrightarrow{E}=(8zx)\cdot\overrightarrow{i}-(3z^2-4x^2)\cdot \overrightarrow{k}" Answer